OpenStudy (anonymous):

lim x --> pi/3 for x (cos x)^2 the answer is zero but how would you solve it with out using a calculator?

OpenStudy (anonymous):

divide all by 2 then multiplybypi sqaured

OpenStudy (anonymous):

wait i think its 4 by using the unit circle whoops

OpenStudy (anonymous):

what i did was pi/3 times cospi/3 squared with x equaling pi/3

OpenStudy (jtvatsim):

Well, if it's the entire cosine squared, then you should see (cos pi/3)^2 = (1/2)^2 = 1/4. Then, multiplying by x = pi/3, is pi/12.